Sep 21, 2017 Calorimeter

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Check in the code, including scripts, that is used for recent calibrations into git (BCAL: Mark)
  2. Add Fcal dead channels into the MC
  3. Mark: Report back on study of eta with in FCAL


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Activities during the down
  4. FCAL Update
  5. BCAL update
  6. Calibrations
  7. Reconstruction and Simulation
  8. NIM articles
  9. Any other business


Attendees: Elton, Mark, Colin (JLab); Will, Mike, Curtis (CMU); Zisis, Andrei, Tegan, Ahmed (UofR); Christina, George (Athens); Matt, Jon (IU)

  1. Announcements
  2. Action Items
    • eta studies in Fcal. Wait till new gain calibrations have been completed
  3. Activities during the down
    1. Fcal (Matt)
      • Power up detector about a month prior to run
      • Jon will extend his time at JLab for the collaboration meeting to work with Colin to bring the detector up
      • Check out dead channels and fix hardware
      • Take some LED runs
      • Redo gain calibrations for the high intensity run for multiple reasons: fix reconstruction, understand resolution and as a basis to adjust HV settings for upcoming run.
    2. Bcal
      • Zisis: Tegan and Ahmed will be at JLab Oct 11-24 and can work to get the Bcal back up and working, take LED runs etc.
      • Elton: We will turn on the Bcal over the next week so that Serguei can work on DAQ and Sasha and work on trigger issues
  4. FCAL Update
  5. BCAL update
  6. Calibrations
    • 2017-01 recon ver02 plans
    • Matt: Concern that time shifts observed in the "calibrated" high intensity 2017 data may be due to some hardware time shifts. Time variation falls on outlines of the DAQ crates, so is suspicious.
    • Colin is working on automating the time constant determination
    • Elton: There will changes to the firmware in the TS, TI and SD to fix some communication issues. These might affect timing problems. If it is completed in time, we may also see changes made to the FADC firmware to address the 4 ns latching offsets we saw in the spring.
    • Mark: Asymmetry in the Bcal timing distribution is likely due to spitoffs and charged tracks that slew the time to later times.
    • The 4 ns secondary peak is likely due to misidentified RF bunches and could be fixed later in the analysis. One needs to understand this and make sure that no events are lost due to cuts which select data from these distributions.
    • Matt: Suggests the use of an exclusive sample, such as omegas, to investigate timing. For example splitoffs could be studies using unused clusters in the exclusive omega sample.
  7. Reconstruction and Simulation
    1. BCAL charged particle clusterizer (Will)
      • Checked changes into the master branch. It is now being used in the latest monitoring run
      • Small changes to code since this was presented at a previous meeting. See July 27 Updates included tuning to optimize parameters.
      • Will: Think about what if any new variables we wish to make available outside the clusterizer (e.g. into the shower or into the REST files).
      • Note that there is a single clusterizer in the code, which contains the "neutral" and "charged" algorithms to be chosen internally.
    2. Neutral showers (Andrei)
      • Working with Irina to use features of clusters to distinguish neutrons from photons. Progress in ongoing and will be summarized in a meeting in the near future.
      • Currently investigating the distribution of the times of contributing cells as a determinant.
    3. MC/Data plots from pi pi events (George)
      • Presented data and MC (pythia) for gp -> pi0pi0p and gp->pi+pi-p samples.
      • Comparisons show reasonable agreement for charged mode, but less so for neutral mode. This is understandable because the dominant charged mode are rhos, which are modeled reasonably well in pythia
      • The ratio neutral/charged ~1-2% for both MC and data. The rho production and Delta++ channels do not have corresponding yields in neutrals.
      • The neutral mode has a rich resonance structure which is missing in pythia.
      • Initial analysis selection without further cuts shows larger differences between data and MC, but these are likely due to different backgrounds and interpretation is unclear unless the selection is better understood.
      • Eugene: Can check pythia output for records of the particles that are being produced to understand what might be expected especially in the neutral mode.
    4. Calorimeter position MC/Sp. 16 Data (Mike)
      • Looking at gp->omega p with omega->pi+pi-pi0 with "loose cuts", but relatively clean sample anyway. 2016 data sample.
      • Finds picket fence structure in phi distribution, both in MC and data (~24 peaks).
      • Will sees 192 peaks in the 2017 data that are due to the phi granularity of the Bcal columns. The 24 peaks might be due to a beating against the histogram bins.
      • Mike also sees a slow -cos(phi) type dependence in the data and a phi dependence at the downstream end of the bcal, which might be more problematic because it is likely to feed into a amplitude analysis. Note: these are not seen in MC.
      • Mark: Note that the data and MC use different constants (i.e. ccdb variations).
      • Matt: Any geometrical features that make the Bcal asymmetric? None that anyone could think of.
      • Mark: Calibration depends on tracking. 2016 data was calibrated before the DC alignment was completed.
      • Elton: Mike should make the same plots for the 2017 data to limit the range of possibilities of this effect.
      • New 2016 data will be reprocessed with all the new improvements once Fcal calibrations are redone.
      • If 2017 data do not show the phi variation, then a recalibration of 2016 data may be necessary. Will notes that the new calibration procedure was not used for 2016 data.
  8. NIM articles (Bcal)
    • Elton has a new version with a large number of comments from Beni
    • We find differences in the constant term in the energy parameterization between Will's pi0 analysis (~1%) and his eta analysis (~3-4%). This is being investigated, but may be due to the dependence of this term on the position along the Bcal. We are trying to understand this effect before putting out a final version of the paper.
    • Zisis: Did Will Brooks respond? Elton: Yes, but does not have time to look at the paper till next week.
    • Elton: Hope to have a final draft available in a week to get feedback from the collaboration.
  9. Any other business